# Redirection for Erasing Memory (REM): Towards a universal unlearning method for corrupted data

### v1

@@ 4 Our Taxonomy of Unlearning Tasks and Limitations of Existing Methods
5 Redirection for Erasing Memory (REM)
We now introduce our core contribution, Redirection for Erasing Memory (REM), the first unlearning method that is performant across our 2D space spanning different regularities and discovery rates, without requiring specialized hyperparameter tuning for each region. The design of REM is based on the following lessons discussed in the previous section: (i) In the setting of partial discovery, finetuning on $\mathcal{D}_{\text{r}}$ causes reinforcing or reintroducing corruption (especially for high-regularity corruptions), but not using $\mathcal{D}_{\text{r}}$ at all would lead to low model utility; (ii) ETD can eliminate the effect of low-regularity corruptions by simply dropping out the memorization partition, but this operation will not remove the effect of high-regularity corruptions, as those are encoded in the generalization partition.
Let $\theta_{o_{1}}=\mathcal{A}(\mathcal{D}_{\text{tr}})$ denote the parameters of the original model, trained using $\mathcal{A}$ on $\mathcal{D}_{\text{tr}}$. Unlike ETD, REM’s choice of $\mathcal{A}$ can be any standard training procedure. REM then applies a post-processing $\mathcal{U}$ on top of $\theta_{o_{1}}$ to unlearn corrupted data by redirecting the effect of corruptions to newly-added parameters $\theta_{o_{2}}$ that are then dropped. We emphasize that, while this design draws inspiration from ETD, it is a fundamentally distinct methodology: while ETD attempts at training time to partition the network into generalization and memorization, REM attempts at post-processing time to partition the network into parameters that are unaffected by the corruptions and ones that capture all corruptions. Furthermore, REM adopts a unique masking strategy to capture high regularity corruptions that cannot be captured by ETD. We detail each REM step from Fig. 2 below and in pseudocode in Alg 1.
@@ 5 Redirection for Erasing Memory (REM)
         2.1 Compute loss $\mathcal{L}_{\text{remove}}(\mathcal{D}_{\text{f}})$ using only the $\theta_{o1}$ part of the model (see Eq. 2)
         2.2 Update $\theta_{o1}$ model part with loss $\mathcal{L}_{\text{step2}}=\mathcal{L}_{\text{remove}}$
@@
     3.2 Compute $\mathcal{L}_{\text{remove}}$ using only the $\theta_{o1}$ part of the model with $\mathcal{D}_{\text{f}}$
     3.3 Update model with the combined loss $\mathcal{L}_{\text{step3}}=\mathcal{L}_{\text{redirect}}-\mathcal{L}_{\text{remove}}$ shown in Eq. 2
@@
Remove corruptions from $\theta_{o_{1}}$ (step 2 of the pseudocode). We achieve this via Negative Preference Optimization (NPO) [30], a SOTA unlearning method in the NLP domain which we translate to our classification problem. NPO is chosen over Potion and Gradient Ascent, the other UL methods available that do not utilize $\mathcal{D}_{\text{r}}$, because Potion destroys model utility to an unrecoverable level at lower regularity tasks (see Fig. 1 (c)) and NPO has demonstrated better performance restoring or healing knowledge, as opposed to removing knowledge in the NLP domain [23], which is closely related to our goal of healing corrupted data. As research progresses, better UL methods that do not utilize $\mathcal{D}_{\text{r}}$ can replace NPO to boost REM performance further. NPO, adapted to classification by using the cross-entropy loss $\mathcal{L}_{\text{CE}}=-\sum_{c=1}^{C}y_{c}\log(\hat{y}_{c})$ , is shown as $\mathcal{L}_{\text{remove}_{\theta_{o_{1}}}}$ in Eq. 2 and used in step 2 of Alg. 1. Note that independence from the retain set during this step is essential to avoid reintroduction of corruptions in $\theta_{o_{1}}$. The downside of this is utility degradation which we address in the next step. The stop condition in Alg. 1 is adapted from [25], which shows that unlearning does not take effect gradually but suddenly as forgetting occurs. The parameter $\gamma$, as shown in [25], is not highly sensitive and should be chosen higher than random chance but lower than model utility.
@@
Repair utility and redirect corruptions into $\theta_{o_{2}}$ instead of relearning in $\theta_{o_{1}}$. At this point, we have removed the effect of corruptions from $\theta_{o_{1}}$ at the expense of utility, via having unlearned on $\theta_{o_{1}}$ using an algorithm that doesn’t use $\mathcal{D}_{\text{r}}$. We now aim to restore utility using $\mathcal{D}_{\text{tr}}$. This operation is risky since, for partial discovery, $\mathcal{D}_{\text{tr}}$ contains discovered and undiscovered corrupted data, in addition to clean data. We therefore design a loss function for this step that can benefit from the clean data of $\mathcal{D}_{\text{tr}}$ while attempting to redirect the corrupted data of $\mathcal{D}_{\text{c}}$ into $\theta_{o_{2}}$, which will then be discarded.
@@
To that end, we design a masking strategy for $\theta_{o_{2}}$ analogously to ETD’s masking strategy in its memorization partition. In particular, every example in $\mathcal{D}_{\text{tr}}$ will receive its own mask in $\theta_{o_{2}}$, which is randomly determined. On the other hand, all examples in $\mathcal{D}_{\text{f}}$ are assigned the same (randomly-determined) mask as each other. Then, directly analogously to ETD, all examples pass through all of $\theta_{o_{1}}$ but each example will only pass through its assigned path in $\theta_{o_{2}}$ based on its mask. The intent is that forcing all forget set examples to share the same mask will cause the resulting path in $\theta_{o_{2}}$ to strongly learn the corruptions, turning into a “channel” for corrupted data. Because we had already unlearned corrupted data from $\theta_{o_{1}}$ in the previous step (a property which we reinforce further by adding NPO to the objective of this step too), we are able to direct the corrupted data of the forget set to this channel in $\theta_{o_{2}}$ as well as possible. Then, undiscovered corrupted examples in the retain set may be partially caught by that channel due to their masks partially overlapping with it. At the same time, clean retain set examples can repair utility, through fine-tuning $\theta_{o_{1}}$ which is kept clean from corruptions due to the previous removing step and the continued application of NPO on $\theta_{o_{1}}$ here too (step 3.2 in pseudocode). The combined loss for this phase is shown below (step 3.3 in pseudocode):
@@
$\mathcal{L}_{\text{step3}}=\underbrace{\frac{2}{\beta}\mathbb{E}\log\sigma\left(-\beta\log\left(\frac{L_{\text{CE}_{\theta_{o_{1}}\cup\theta_{o_{2}}}}(\mathcal{D}_{\text{tr}})}{L_{\text{CE}_{ref}}(\mathcal{D}_{\text{tr}})}\right)\right)}_{\text{$\mathcal{L}_{\text{redirect}_{\theta_{o_{1}}\cup\theta_{o_{2}}}}$}}-\underbrace{\frac{2}{\beta}\mathbb{E}\log\sigma\left(-\beta\log\left(\frac{\mathcal{L}_{\text{CE}_{\theta{o1}}}(\mathcal{D}_{\text{f}})}{L_{\text{CE}_{ref}}(\mathcal{D}_{\text{f}})}\right)\right)}_{\text{$\mathcal{L}_{\text{remove}_{\theta_{o_{1}}}}$}}.$ (2)
@@
The reference models here (“ref”) correspond to the initialization of the respective weights (both $\theta_{o_{1}}$ and $\theta_{o_{2}}$ in the first term; only $\theta_{o_{1}}$ in the second). This formulation is similar to Direct Preference Optimization (DPO) [22] with the key difference that we perform each part on a different model (i.e. with a different set of parameters active and a different forward pass that is example-dependent).
@@ 6 Results & Discussion
Figure 3: Comparison of UL methods on two model capacity levels using ResNet-9 & CIFAR10 with 1000 corrupted samples, three regularity levels and 10 discovery rates (10%-100%). REM (IDEAL) provides and upper limit with perfect knowledge of manipulated samples for mask assignment. ETD and REM are not reported for 100% capacity as reserve capacity is needed for $\theta_{o_{2}}$ of the model. Error bars reflect ±$\pm 1$ SEM.
@@
Metrics. “Healed” computes the clean (i.e. uncorrupted label) accuracy on the corrupted $\mathcal{D}_{\text{f}}$ inputs. Utility computes the accuracy on a clean test set without any corruptions. Utility $\times$ Healed provides a multiplicative score, where higher is better, to ensure that the method not only removes the corrupted data but also retains model utility. Figs. 1 (b-g) are shaded using this aggregate metric. Reported errors reflect ±$\pm 1$ standard error of the mean (SEM).
Method comparison. Fig. 1 shows that REM is the only method that performs strongly on all regions of our 2D space. We additionally provide aggregate numerical scores in Fig. 3, that are computed as the average of the Utility $\times$ Healed metric across runs for all discovery rates and regularities. We make the following observations. For all methods, the score increase for the 50% capacity model over the 100% model is due to less memorization of lower regularity corruptions due to less overparameterization. We observe that ETD significantly outperforms the approach of simply training a smaller original model (“no UL” in Fig. 3).
In fact, we find that ETD alone is a strong baseline that was not previously considered. It outperforms all prior unlearning methods in terms of the aggregate scores, due to the effectiveness of memorization mitigation on lower regularity tasks. Further, as also discussed by [9], retraining a model from scratch is not the gold standard for partial discovery tasks as the undiscovered corruption is reintroduced into the model, leading to poor performance. Potion does not perform well on this aggregate metric due to failure on the lower regularity tasks. A surprising discovery is that the simple baseline of Gradient Ascent beats many other methods on our 2D space in terms of aggregate score. This is a surprising as this baseline was believed to be weak and was even omitted from prior comparisons [9]. However, we note that Ascent fails across entire slices of the 2D space as shown in Fig. 1, which is hidden in the aggregation but may be problematic in practice. REM outperforms all prior metrics and even comes close to REM (IDEAL), an upper bound that has knowledge of all corrupted data when forming masks at unlearning time.
Table 1: Aggregate results of methods on CIFAR10 with ResNet-9 and ViT with SVHN (50% capacity each) and 1000 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed (ResNet-9). Potion ViT experiment failed due to OOM (Out of memory) with an A100 (40GB VRAM). Error reflects ±$\pm 1$ SEM.
@@
ETD UL type Healed Utility Utility×$\times$×Healed Utility×$\times$×Healed
@@
Unlearning on an ETD model. Tab. 1 shows that REM is the best performing method when used either with or without ETD training. The choice of whether to use ETD pretraining leads to a trade-off, as shown in Fig. 4: ETD improves performance on the lower regularity tasks but its training procedure causes a model utility deficit as the model used during training differs from the model at inference time. Other unlearning methods show an increased combined score of Utility $\times$ Healed when combined with ETD but do not come close to REM. This is due to the shortcomings of the last used method affecting the results. For example, (i) in SCRUB, the reintroduction of the poison trigger in partial discovery settings causes failure in high regularity tasks, (ii) Potion causes significant model utility damage in the lower regularity settings negating the benefit of ETD in this space.
Prevention versus postprocessing. As shown in Fig. 3, less overparameterization (smaller capacity) is effective in partly mitigating memorization of corruptions during training. We report extreme cases of this in the appendix (see Fig. 7), showing that by restricting model capacity to very low levels, no space for memorization remains, effectively solving lower regularity tasks but at the high cost of model utility. Therefore other mitigation actions during training, such as ETD for instance, should be preferred over extreme capacity restriction. Based on these insights, we argue that high regularity corruptions are the hardest to address during training. We suggest that this also applies to other domains such as NLP, where verbatim memorization can be addressed in simple ways such as with the Goldfish loss [11] but concepts (high regularity) are difficult to mitigate during training.
Table 2: REM ablation using CIFAR10 with ResNet-9 analogous to Tab. 1. Redirection (3.1) using $\mathcal{D}_{\text{tr}}$ instead of just finetuning on $\mathcal{D}_{\text{r}}$ is essential while step 3.2 only matters when the model is not trained with ETD. No use of 3.1 & 3.2 simplifies to NPO, as $\theta_{o_{2}}$ is unused. Error reflects ±$\pm 1$ SEM.

### v2

@@ 5 Redirection for Erasing Memory (REM)
Remove corruptions from $\theta_{o_{1}}$ (step 2 of the pseudocode). We achieve this via Negative Preference Optimization (NPO) (Zhang et al., 2024a), a SOTA unlearning method in the NLP domain which we translate to our classification problem. NPO is chosen over Potion and Gradient Ascent, the other UL methods available that do not utilize $\mathcal{D}_{\text{r}}$, because Potion destroys model utility to an unrecoverable level at lower regularity tasks (see Fig. 1 (c)) and NPO has demonstrated better performance restoring or healing knowledge, as opposed to removing knowledge in the NLP domain (Rezaei et al., 2025), which is closely related to our goal of healing corrupted data. As research progresses, better UL methods that do not utilize $\mathcal{D}_{\text{r}}$ can replace NPO to boost REM performance further. NPO, adapted to classification by using the cross-entropy loss $\mathcal{L}_{\text{CE}}=-\sum_{c=1}^{C}y_{c}\log(\hat{y}_{c})$ , is shown as $\mathcal{L}_{\text{remove}_{\theta_{o_{1}}}}$ in Eq. 2 and used in step 2 of Alg. 1. Note that independence from the retain set during this step is essential to avoid reintroduction of corruptions in $\theta_{o_{1}}$. The downside of this is utility degradation which we address in the next step. The stop condition in Alg. 1 is adapted from Schoepf et al. (2024b), which shows that unlearning does not take effect gradually but suddenly as forgetting occurs. The parameter $\gamma$, as shown in Schoepf et al. (2024b), is not highly sensitive and should be chosen higher than random chance but lower than model utility.
@@
   2.1 Compute loss $\mathcal{L}_{\text{remove}_{\theta_{o_{1}}}}(\mathcal{D}_{\text{f}})$ using only the $\theta_{o1}$ part of the model (see Eq. 2)
   2.2 Update $\theta_{o1}$ model part with loss $\mathcal{L}_{\text{step2}}={\color[rgb]{0,0,0}-}\mathcal{L}_{\text{remove}_{\theta_{o_{1}}}}$
@@
  3.2 Compute $\mathcal{L}_{\text{remove}_{\theta_{o_{1}}}}(\mathcal{D}_{\text{f}})$ using only the $\theta_{o1}$ part of the model (see $\mathcal{L}_{\text{remove}_{\theta_{o_{1}}}}$ in Eq. 2)
  3.3 Update model with the combined loss $\mathcal{L}_{\text{step3}}=\mathcal{L}_{\text{redirect}_{\theta_{o_{1}}\cup\theta_{o_{2}}}}-\mathcal{L}_{\text{remove}_{\theta_{o_{1}}}}$ shown in Eq. 2
@@
Repair utility and redirect corruptions into $\theta_{o_{2}}$ instead of relearning in $\theta_{o_{1}}$. At this point, we have removed the effect of corruptions from $\theta_{o_{1}}$ at the expense of utility, via having unlearned on $\theta_{o_{1}}$ using an algorithm that doesn’t use $\mathcal{D}_{\text{r}}$. We now aim to restore utility using $\mathcal{D}_{\text{tr}}$. This operation is risky since, for partial discovery, $\mathcal{D}_{\text{tr}}$ contains undiscovered corrupted data. We therefore design a loss function for this step that can benefit from the clean data of $\mathcal{D}_{\text{tr}}$ while attempting to redirect the corrupted data of $\mathcal{D}_{\text{c}}$ into $\theta_{o_{2}}$, which will then be discarded.
@@
To that end, we design a masking strategy for $\theta_{o_{2}}$ analogously to ETD’s masking strategy in its memorization partition. In particular, every example in $\mathcal{D}_{\text{tr}}$ will receive its own randomly determined mask in $\theta_{o_{2}}$. All examples in $\mathcal{D}_{\text{f}}$ are assigned the same (randomly-determined) mask as each other. Then, directly analogously to ETD, all examples pass through all of $\theta_{o_{1}}$ but each example will only pass through its assigned path in $\theta_{o_{2}}$ based on its mask. The intent is that forcing all forget set examples to share the same mask will cause the resulting path in $\theta_{o_{2}}$ to strongly learn the corruptions, turning into a “channel” for corrupted data. Utilizing the same mask for all corrupted samples intuitively seems like a problem, as the model may learn generalized knowledge in the added capacity when a neuron is present in multiple masks of samples that share generalizable knowledge. Specifically, when the added capacity is dropped, this generalized knowledge will be lost and model utility drops. This can indeed be true if the masking strategy is active from the start of training. However, in the post-hoc unlearning setting of REM, the model is already trained. Existing generalization will not switch over to the added capacity mask, as the path of least resistance is to reuse the already present neurons in the model that contain this information. We therefore first need to remove the existing information from these neurons (the Remove step of REM) to be able to redirect them. Because we had already unlearned corrupted data from $\theta_{o_{1}}$ in the previous step (a property which we reinforce further by adding NPO to the objective of this step), we are able to direct the corrupted data of the forget set to this channel in $\theta_{o_{2}}$.
At the same time as the redirection of corruptions, clean retain set examples can repair utility through fine-tuning $\theta_{o_{1}}$ which is kept clean from corruptions due to the previous removing step and the continued application of NPO on $\theta_{o_{1}}$ (Alg. 1 step 3.2). The combined loss (Alg. 1 step 3.3) is shown with $\sigma$ as the sigmoid function:
@@
$\mathcal{L}_{\text{step3}}=\underbrace{\frac{2}{\beta}\mathbb{E}\log\sigma\left(-\beta\log\left(\frac{{\color[rgb]{0,0,0}\mathcal{L}}_{\text{CE}_{\theta_{o_{1}}\cup\theta_{o_{2}}}}(\mathcal{D}_{\text{tr}})}{{\color[rgb]{0,0,0}\mathcal{L}}_{\text{CE}_{ref}}(\mathcal{D}_{\text{tr}})}\right)\right)}_{\text{$\mathcal{L}_{\text{redirect}_{\theta_{o_{1}}\cup\theta_{o_{2}}}}$}}-\underbrace{\frac{2}{\beta}\mathbb{E}\log\sigma\left(-\beta\log\left(\frac{{\color[rgb]{0,0,0}\mathcal{L}}_{\text{CE}_{\theta{o1}}}(\mathcal{D}_{\text{f}})}{{\color[rgb]{0,0,0}\mathcal{L}}_{\text{CE}_{ref}}(\mathcal{D}_{\text{f}})}\right)\right)}_{\text{$\mathcal{L}_{\text{remove}_{\theta_{o_{1}}}}$}}.$ (2)
@@
The reference models (“ref”) correspond to the initialization of the respective weights (both $\theta_{o_{1}}$ and $\theta_{o_{2}}$ in the first term; only $\theta_{o_{1}}$ in the second). This formulation is similar to Direct Preference Optimization (DPO) (Rafailov et al., 2023) with the key difference that we perform each part on a different model (i.e. a different set of parameters active and an example-dependent forward pass).
@@ 6 Results & Discussion
Figure 3: Comparison of UL methods on two model capacity levels using ResNet-9 & CIFAR10 with 1000 corrupted samples, three regularity levels and 10 discovery rates (10%-100%). REM (IDEAL) provides and upper limit with perfect knowledge of manipulated samples for mask assignment. ETD and REM are not reported for 100% capacity as reserve capacity is needed for $\theta_{o_{2}}$ of the model. Error bars reflect $\pm 1$ SEM.
@@
Metrics. “Healed” computes the clean (i.e. uncorrupted label) accuracy on the corrupted $\mathcal{D}_{\text{c}}$ inputs. Utility computes the accuracy on a clean test set without any corruptions. Utility $\times$ Healed provides a multiplicative score, where higher is better, to ensure that the method not only removes the corrupted data but also retains model utility, analogous to Zhao et al. (2024). Figs. 1 (b-g) are shaded using this aggregate metric. Reported errors reflect $\pm 1$ standard error of the mean (SEM).
Table 1: Aggregate results of methods on CIFAR10 with ResNet-9 and ViT with SVHN (50% capacity each) and 1000 corrupted samples with 10 discovery rate levels (10%-100%) and 3 regularities. Presented in descending order of Utility $\times$ Healed (ResNet-9). Potion ViT experiment failed due to OOM (Out of memory) with an A100 (40GB VRAM). Error reflects $\pm 1$ SEM.
@@
[…]
@@
Method comparison. Fig. 1 shows that REM is the only method that performs strongly on all regions of our 2D space. We additionally provide aggregate numerical scores in Fig. 3, that are computed as the average of the Utility $\times$ Healed metric across runs for all discovery rates and regularities. We make the following observations. For all methods, the score increase for the 50% capacity model over the 100% model is due to less memorization of lower regularity corruptions due to less overparameterization. We observe that ETD significantly outperforms the approach of simply training a smaller original model (“no UL” in Fig. 3). In fact, we find that ETD alone is a strong baseline that was not previously considered. It outperforms prior unlearning methods in terms of the aggregate scores, due to the effectiveness of memorization mitigation on lower regularity tasks. Further, as also discussed by Goel et al. (2024), retraining a model from scratch is not the gold standard for partial discovery tasks as the undiscovered corruption is reintroduced into the model, leading to poor performance. Potion does not perform well on this aggregate metric due to failure on the lower regularity tasks. A surprising discovery is that the simple baseline of Gradient Ascent beats many other methods on our 2D space in terms of aggregate score. This is a surprising as this baseline was believed to be weak and was even omitted from prior comparisons (Goel et al., 2024). However, we note that Ascent fails across entire slices of the 2D space as shown in Fig. 1, which is hidden in the aggregation but may be problematic in practice. REM outperforms all prior metrics and even comes close to REM (IDEAL), an upper bound that has knowledge of all corrupted data when forming masks at unlearning time.
@@
Unlearning on an ETD model. Tab. 1 shows that REM is the best performing method when used either with or without ETD training. The choice of whether to use ETD pretraining leads to a trade-off, as shown in Fig. 4: ETD improves performance on the lower regularity tasks but its training procedure causes a model utility deficit as the model used during training differs from the model at inference time. Other unlearning methods show an increased combined score of Utility $\times$ Healed when combined with ETD but do not come close to REM. This is due to the shortcomings of the last used method affecting the results. For example, (i) in SCRUB, the reintroduction of the poison trigger in partial discovery settings causes failure in high regularity tasks, (ii) Potion causes significant model utility damage in the lower regularity settings negating the benefit of ETD.
Prevention versus postprocessing. As shown in Fig. 3, less overparameterization (smaller capacity) is effective in partly mitigating memorization of corruptions during training. We report extreme cases of this in the appendix (see Fig. 8), showing that by restricting model capacity to very low levels, no space for memorization remains, effectively solving lower regularity tasks but at the high cost of model utility. Therefore other mitigation actions during training, such as ETD for instance, should be preferred over extreme capacity restriction. Based on these insights, we argue that high regularity corruptions are the hardest to address during training. We suggest that this also applies to other domains such as NLP, where verbatim memorization can be addressed in simple ways such as with the Goldfish loss (Hans et al., 2024) but concepts (high regularity) are difficult to mitigate.
Redirection of corruptions. We show the effectiveness of our redirection strategy using ResNet9, 50% discovery rate, and the poison trigger task over 5 unlearning epochs. As shown in Fig. 5, the base model starts off with 99.0% accuracy on the corrupted data (i.e. nearly every sample adversely affected). During unlearning, the base model (i.e. with no added parameters active) accuracy on corrupted data drops to around 10%, which is random chance with 10 classes. This is made possible because at the same time, the corruptions were rerouted into the newly added parameter partition, as shown by the accuracy on the corrupted data for the base model plus the active parameters of the expansion. These results clearly show that the corrupted examples are indeed rerouted, enabling unlearning without the problem of reintroduction that makes prior methods fail.
Table 2: REM ablation using CIFAR10 with ResNet-9 analogous to Tab. 1. Redirection (3.1) using $\mathcal{D}_{\text{tr}}$ instead of just finetuning on $\mathcal{D}_{\text{r}}$ is essential while step 3.2 only matters when the model is not trained with ETD. No use of 3.1 & 3.2 simplifies to NPO, as $\theta_{o_{2}}$ is unused. Error reflects $\pm 1$ SEM.
@@
Figure 5: Accuracy on corrupted data for the base model (i.e. the existing trained model) and the base model plus the additional active parameters from the added mask. Epoch 0 is the starting point where no unlearning has taken place. As unlearning takes place, the corruptions are redirected into the additional capacity of the model and removed from the base model. This enables the base model to be free of corruptions by dropping the added capacity after unlearning.